Please help place the following items into the transactions!i More Info The Baxter Company was incorporated on September 1, 20X0. Baxter had 20 holders of common stock. Reya Baxter, the
Answer #1
accounting equation:
Assets = Liabilities + stockholders' Equity
S.No. Cash+ AR + Inv.+ Prep Rent+ Equip&Fixt. = NP + AP + Paid in Capital + Retained Earnings
a 600000 = 600000
b -25000 25000 =
c 65000 = 65000
d1 28000 82000 = 110000
d2 -31000 = -31000
e 5000 -5000 =
f -45000 = -45000
g -18000 115200 = 97200
h -4500 4500 =
i -3300 = -3300 (110000*3%)
j -31000 = -31000
k -1600 = -1600
l                                                          -1500                       =                                                       -1500
Balance 506200 77000 59000 3000 113600 = 97200 20000 600000 41600
G. total 758800 = 758800